OpenStudy (adamaero):

Help putting two cross products into matrices... MEDAL (-i + j)X(F_A)k & 2jX[(E_x)i + (E_z)k]

OpenStudy (adamaero):

\[(-i + j)X(F_{A}k) + (2j)X(E_{x}i + E_{z}k)\]

zepdrix (zepdrix):

I guess I'm thinking of it more like this:\[\Large\rm \left<-1,1,0\right>\times\left<0,0,F_A\right>+\left<0,2,0\right>\times\left<E_x,0,E_z\right>\]Mmmm ya? Did I set that up correctly?

zepdrix (zepdrix):

It just seems like it would be easier to work with from that form.

OpenStudy (adamaero):